Find the capacitance of a spherical capacitor filled with a dielectric of dielectric 

constant K.


1
Expert's answer
2021-10-27T10:14:04-0400

"potential \\V_1=\frac{k(-q)}{Kr_2}"

Potential plate second


"V_2=\frac{k(+q)}{Kr_1}"

We know that

"\u2206V=V_1-V_2"

Put value

"∆V=\frac{q}{4\pi\epsilon_0K}[\frac{1}{r_1}-\frac{1}{r_2}]"


"∆V=\frac{q}{4\pi\epsilon_0K}[\frac{r_1-r_2}{r_1r_2}]"

"∆V=\frac{q}{4\pi\epsilon_0K}\frac{r_1-r_2}{r_1r_2}"

We know that capacitance

"C_0=\frac{q}{∆V}"

Put value

"C_0=4\pi\epsilon_0K\frac{r_1r_2}{r_1-r_2}"